In a natural gas pipeline, the velocity v (in m/s) of the gas as a function of the distance x (in cm) from the wall of the pipe is given by v = 5.2x - x^2 . Determine x for v = 4.8 m/s.
Answer by amarjeeth123(569)

For v=4.8 m/s 4.8=5.2x-x^2
x^2-5.2x+4.8=0
x^2-4x-1.2x+4.8=0
x(x-4)-1.2(x-4)=0
(x-1.2)(x-4)=0
x=1.2 or x=4.